# Linguistically Grounded Analysis of Language Models using Shapley Head Values

### v1

@@ 1 Introduction
    We release an implementation of our SHV-based probing and pruning methods.111 Repository under preparation, and can be shared upon request.
@@ 2 Background
In this section, we describe previous work in probing, aiming to discern the importance of individual training data points or language model components in the processing of specific phenomena – including language- and task-specificity, morphosyntactic phenomena, and bias (Section 2.1). Then, we introduce Shapley Values, and their role in analysing the contribution of attention heads inside a language model towards a target task (Section 2.2). Finally, we explain pruning, a method of switching off neurons or – in the case of structural pruning –, attention heads using binary masks (Section 2.3).
@@ 2.1 Probing
Previous work in attribution focuses on identifying the role that either training features or model components play with respect to various phenomena. Training data attribution is used to investigate, e.g., to what extent language-specific sub-networks rely on in-language examples from the training data when making predictions (Choenni et al., 2023). Foroutan et al. (2022) uses iterative magnitude pruning (Frankle and Carbin, 2019) to isolate sub-networks specific to languages and tasks like masked language modelling, named entity recognition, and natural language inference. Structured pruning is used for identifying attention heads important for dialogue summarisation (Liu and Chen, 2023) and cross-lingual natural language inference, e.g., with Shapley Head Values, (Held and Yang, 2023). Extrinsic probing efforts such as from Ács et al. (2023) target whether morphological information is encoded by language models. On the other hand, work in intrinsic probing aims to reveal how exactly linguistic information is structured within a model (Dalvi et al., 2019; Torroba Hennigen et al., 2020; Stanczak et al., 2022)
Another popular technique is causal mediation analysis (CMA), a causal method of identifying the importance of a model component to a target phenomenon (Pearl, 2001). Under CMA, two effects on the prediction of a model are compared: a direct effect of input intervention – such as a text edit – and an indirect effect. The relative degree of the indirect effect compared to the direct effect reveals the significance of the target component to the target phenomenon. CMA is used to investigate the role of individual neurons and attention heads in mediating gender bias (Vig et al., 2020), and to isolate neurons responsible for subject-verb agreement in English (Finlayson et al., 2021), and across a number of other languages (Mueller et al., 2022). A limitation of CMA, however, is that text edit operations are necessarily word-level to isolate other factors such as sensitivity to specific syntactic features. This makes it impossible to accommodate a number of morphosyntactic constructions such as island effects (see Table 1). Instead, we use SHVs which are particularly well-suited to probe for subnetworks responsible for processing of morphosyntactic phenomena.
@@ 3.1 Deriving SHVs
Following Held and Yang (2023), we define SHV $\varphi_{h}$, for a single attention head $\textrm{Att}_{h}\in A$, to represent the mean performance improvement on the characteristic function $V$ – derived from the evaluation metric $M_{v}$ and described in Section 3.2.2 – if $\textrm{Att}_{h}$ contributes to the inference. To be able to remove or add attention heads at will, we augment our target models with a gate $G_{h}=\{0,1\}$ for each attention head. When $G_{h}=0$, the head $\textrm{Att}_{h}$ is removed from the inference and does not contribute to the output of the transformer. The derivation of $\varphi_{h}$ requires the contribution of the head $\textrm{Att}_{h}$ to be measured across all $Q$ permutations of the other gates, see Equation 1.
@@ 3.2.2 Grammaticality Judgement Task
We define $M_{v}$ mentioned in Section 3.1 for the SHV derivation as accuracy on a custom grammaticality judgement task: a classifier has to output a binary label signifying which element of a sentence pair $(s_{1},s_{2})$ is grammatical. The same metric is used to derive pruning performance. Both sentences are drawn from the same BLiMP paradigm $p$, member of the set of all paradigms $\mathcal{P}$, $p$ shuffled prior to the selection of $(s_{1},s_{2})$. Shuffling is done to ensure that the classifier focuses on the underlying morphosyntactic phenomenon represented (or violated) in the sentences, rather than the surface features of $(s_{1},s_{2})$.
The performance of the classifier is measured in terms of accuracy, and it has to assign label 00 or $1$, depending on if $s_{1}$ or $s_{2}$ is grammatical:
@@
The advantage of this novel task formulation is that – unlike previously suggested approaches such as simple text edits – it can flexibly incorporate morphosyntactic phenomena where the grammatical and ungrammatical sentences differ in more than a single word (see e.g., Ellipsis or Island effects in Table 1).
@@ 3.2.3 Fine-Tuning Setup
We carry out our experiments and our analysis using the monolingual English models BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019). BERT has 110 million parameters and was trained on 16GB data, while RoBERTa has 125 million parameters and was trained on 160GB. We follow Held and Yang (2023) in deriving SHVs (see Section 3.1). However, rather than fully fine-tuning our target model weights for the target task, we use low-rank adapters (LoRA; Hu et al. 2021) implemented with the PEFT library. [^3] See https://huggingface.co/docs/peft/index. The main advantages of LoRA modules are their speed of training, and, more importantly, that they do not modify the original model weights. This has the added advantage that it enables us to isolate and evaluate the pretrained linguistic knowledge of the models we consider.
@@ 3.3 Interpretation
Figure 3: Mean $\Delta$ accuracy values when pruning the top $n$ heads across paradigms.
@@ 4 Results
Table 2: BLiMP paradigms in the relevant BERT ($\blacklozenge$) and RoBERTa ($\blacklozenge$) clusters. Most paradigms appear in the clusters of both models, these are marked with $\lozenge$. There are two different Binding clusters in BERT, the paradigms in the smaller cluster are underlined.
@@ 5 Discussion
In the following, we discuss our most important findings regarding the clusters created via SHV attributions. The membership of these clusters reflect which morphosyntactic phenomena language models like BERT and RoBERTa generalise over via using the same subnetworks.
@@
The difference between in-cluster and out-of-cluster pruning impact on $\Delta$ accuracy indicates we identified the majority of attention heads representing the full construction-specific subnetwork (see Figure 5). In a subset of cases this is not the case: either the subnetwork is larger, or the relevant knowledge is more widely distributed across model components.
@@ 6 Conclusion
A major limitation for our work is the difficulty of scaling it to new languages and to new language models. First, the coverage of datasets of linguistic minimal pairs for languages other than English lags behind that of BLiMP. Next, linguistically grounded clustering analysis requires an in-depth understanding of the specific morphosyntactic phenomena in a language, as well as considerable manual effort. This effort is a bottleneck on how readily the analysis can be carried out for other language models and, of course, other languages. Additionally, sentences in the BLiMP dataset are generated from expert templates using a predefined vocabulary, which means they tend to be semantically empty as well as heavily formulaic. However, since this is true for both grammatical and ungrammatical sentences, this fact is not likely to cause problems for our attributions. Furthermore, the derivation of SHVs requires the use of sampling techniques and truncation heuristics to make computation tractable. These techniques introduce uncertainty, which means there is a slight chance they might influence the accuracy of the attribution process. Finally, we are carrying out attributions on the level of the attention head, rather than on the level of the neuron. This improves tractability and facilitates the qualitative analysis, but it may result in a loss of granularity. In future work, it might be worth exploring what we can gain from neuron-level attributions.

### v2

@@ 1 Introduction
    We release an implementation of our SHV-based probing and pruning methods.
@@ 2.1 Probing
Previous work in attribution focuses on identifying the role that either training features or model components play with respect to various phenomena. Training data attribution is used to investigate, e.g., to what extent language-specific sub-networks rely on in-language examples from the training data when making predictions (Choenni et al., 2023). Foroutan et al. (2022) uses iterative magnitude pruning (Frankle and Carbin, 2019) to isolate sub-networks specific to languages and tasks like masked language modelling, named entity recognition, and natural language inference. Structured pruning is used for identifying attention heads important for dialogue summarisation (Liu and Chen, 2023) and cross-lingual natural language inference, e.g., with Shapley Head Values, (Held and Yang, 2023). Extrinsic probing efforts such as from Ács et al. (2023) target whether morphological information is encoded by language models. On the other hand, work in intrinsic probing aims to reveal how exactly linguistic information is structured within a model (Dalvi et al., 2019; Torroba Hennigen et al., 2020; Stanczak et al., 2022).
Another popular technique is causal mediation analysis (CMA), a causal method of identifying the importance of a model component to a target phenomenon (Pearl, 2001). Under CMA, two effects on the prediction of a model are compared: a direct effect of input intervention – such as a text edit – and an indirect effect. The relative degree of the indirect effect compared to the direct effect reveals the significance of the target component to the target phenomenon. CMA is used to investigate the role of individual neurons and attention heads in mediating gender bias (Vig et al., 2020), and to isolate neurons responsible for subject-verb agreement in English (Finlayson et al., 2021), and across a number of other languages (Mueller et al., 2022). A limitation of CMA, however, is that text edit operations are necessarily word-level to isolate other factors such as sensitivity to specific syntactic features. This makes it challenging to accommodate a number of morphosyntactic constructions such as island effects (see Table 1). Instead, we use SHVs which are particularly well-suited to probe for subnetworks responsible for processing of morphosyntactic phenomena.
Ours is certainly not the first paper to carry out the analysis of language model skills using linguistic grounding, but we believe our work is differentiated by the wider coverage in terms of morphosyntactic phenomena, as well as our efforts to localise linguistic knowledge. Previous work explores the relation between self-attention of input tokens and dependency links (Htut et al., 2019; Clark et al., 2019). Linzen and Baroni (2021) analyse the capabilities of LSTM and GRU models on long-distance dependencies, while Wilcox et al. (2024) use GPT-2 and GPT-3 to look at long-distance dependencies and island constraints.
@@ 3.1 Deriving SHVs
Following Held and Yang (2023), we define SHV $\varphi_{h}$, for a single attention head $\textrm{Att}_{h}\in A$ [^1] Representing the set of all attention heads., to represent the mean performance improvement on the characteristic function $V$ – as derived from performance on the evaluation metric described in Section 3.2.2 – if $\textrm{Att}_{h}$ contributes to the inference. To be able to remove or add attention heads at will, we augment our target models with a gate $G_{h}=\{0,1\}$ for each attention head. When $G_{h}=0$, the head $\textrm{Att}_{h}$ is removed from the inference and does not contribute to the output of the transformer. The derivation of $\varphi_{h}$ requires the contribution of the head $\textrm{Att}_{h}$ to be measured across all $Q$ permutations of the other gates, see Equation 1.
@@ 3.2.2 Grammaticality Judgement Task
We define the evaluation metric mentioned in Section 3.1 for the SHV derivation as accuracy on a custom grammaticality judgement task: a classifier has to output a binary label signifying which element of a sentence pair $(s_{1},s_{2})$ is grammatical. The same metric is used to derive pruning performance. Both sentences are drawn from the same BLiMP paradigm $p$, member of the set of all paradigms $\mathcal{P}$, $p$ shuffled prior to the selection of $(s_{1},s_{2})$. Shuffling is done to facilitate the classifier focusing on the underlying morphosyntactic phenomenon represented – or violated – in the sentences, as opposed to merely the surface features of $(s_{1},s_{2})$.
The performance of the classifier is measured in terms of accuracy, and it assigns label 00 or $1$, depending on if $s_{1}$ or $s_{2}$ is grammatical:
@@
The advantage of this novel task formulation is that – unlike other approaches such as simple text edits – it can flexibly incorporate morphosyntactic phenomena where the grammatical and ungrammatical sentences differ in more than a single word (see e.g., Ellipsis or Island effects in Table 1).
@@ 3.2.3 Fine-Tuning Setup
We carry out our experiments and our analysis using the monolingual English models BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019). BERT has 110 million parameters and was trained on 16GB of data, while RoBERTa has 125 million parameters and was trained on 160GB. We follow Held and Yang (2023) in deriving SHVs (see Section 3.1). However, rather than fully fine-tuning our target model weights for the target task, we use low-rank adapters (LoRA; Hu et al. 2021) implemented with the PEFT library. [^3] See https://huggingface.co/docs/peft/index. The main advantages of LoRA modules are their speed of training, and, more importantly, that they merely emphasise information that is already present in the original weights as opposed to reshaping the original model weights (Hu et al., 2021). This has the added advantage that it enables us to isolate and evaluate the pretrained linguistic knowledge of the models we consider.
@@ 3.3 Interpretation
Figure 3: Mean $\Delta$ accuracy values drop in a near-linear fashion when pruning up to the top $n$ heads across paradigms
@@ 4 Results
Table 2: Clusters that emerge in both models with shared paradigms marked with $\lozenge$. A subset of paradigms only appear in in BERT ($\blacklozenge$) or RoBERTa ($\blacklozenge$). Clusters are named by the authors based on majority membership. [^5] There are two different Binding clusters in BERT, the paradigms in the smaller cluster are underlined.
@@ 5 Discussion
In the following, we discuss our most important findings regarding the clusters created via SHV attributions. The membership of these clusters indicates which morphosyntactic phenomena are processed using the same subnetworks by BERT and RoBERTa, thus showing how the models generalise over these constructions.
@@
The difference between in-cluster and out-of-cluster pruning impact on $\Delta$ accuracy indicates we identified the majority of attention heads representing the full construction-specific subnetwork (see Figure 5). In a subset of cases this is not the case: either the subnetwork is larger, or the relevant knowledge is more widely distributed across model components. Three out of the six clusters show a significant difference between in-cluster and out-of-cluster $\Delta$ accuracy. This is very unlikely to occur randomly: in our experiments with 125 random clusters of varying sizes, we found it to happen only in two cases.
@@ 6 Conclusion
A major limitation for our work is the difficulty of scaling it to new languages and to new language models. First, the coverage of datasets of linguistic minimal pairs for languages other than English lags behind that of BLiMP. Next, linguistically grounded clustering analysis requires an in-depth understanding of the specific morphosyntactic phenomena in a language, as well as considerable manual effort. This effort is a bottleneck on how readily the analysis can be carried out for other language models and, of course, other languages. Additionally, sentences in the BLiMP dataset are generated from expert templates using a predefined vocabulary, which means they tend to be semantically empty as well as heavily formulaic. However, since this is true for both grammatical and ungrammatical sentences, this fact is not likely to cause problems for our attributions. Furthermore, the derivation of SHVs requires the use of sampling techniques and truncation heuristics to make computation tractable. These techniques introduce uncertainty, which means there is a slight chance they might influence the accuracy of the attribution process. Moreover, the binary gates used to zero attention head activations may harm model performance since these zero-values are out-of-distribution for the language model. This means the actual importance of attention heads somewhat difficult to discombobulate from the impact of the zero ablations. Finally, we are carrying out attributions on the level of the attention head, rather than on the level of the neuron. This improves tractability and facilitates the qualitative analysis, but it may result in a loss of granularity. In future work, it might be worth exploring what we can gain from neuron-level attributions.
@@
Acknowledgements
MF and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454).